I am 62 and am tutoring a 50-year old woman who needs to pass the entrance exam for a college algebra class at the University of Maine in Augusta this Fall. I have done this question by calulator for an answer of B=1.08, but she reduced the fractions longhand and got B=0.1. Doing it her way step-by-step I see the problem is in the cross-products, but I cannot write it out in a step-by-step fashion to get the right answer either. Would you please write out the step-by-step reduction for me? The question is:
3¼/2 = B/two-thirds, that is 3 and one-quarter over 2 = B over two-thirds.
Thank you. I see my student again next Tuesday, June 5th, Bob Gross

Found 2 solutions by Theo, lwsshak3:
Answer by Theo(13342)

your equation is:
(3+1/4) / 2 = B / (2/3)
you can make an improper fraction out of 3+1/4.
it becomes (13/4)
your equation becomes:
(13/4) / 2 = B / (2/3)
cross multiply to get:
(13/4) * (2/3) = 2 * B
this is equivalent to:
(13*2) / (4*3) = 2 * B
simplify this further to get:
26 / 12 = 2 * B
divide both sides of this equation by 2 to get:
13 / 12 = B
decimal equivalent of 13 / 12 = 1.08333333333 which rounds to 1.08 or 1.083 or whatever detail you want to go out to.

Answer by lwsshak3(11628)

3¼/2 = B/two-thirds
3 and 1/4=13/4
rewrite equation:
(13/4)/2=B/(2/3)
cross-multiply
2B=(2/3)(13/4=26/12
divide by 2
B=26/24=13/12
If an exact answer is wanted, you must leave it in this fraction form.
If a decimal answer is wanted, you will get 13/12=1.08333...
so, you would probably round it off depending on how many decimal places are wanted.
one decimal place=1.1
two decimal places=1.08
three decimal places=1.083
etc.
hope this helps